OpenStudy (anonymous):

Create an equation with point-slope form for: passes through (1, 2) with slope of -5/3

OpenStudy (anonymous):

Point slope form is (y - y1) = m(x - x1)

OpenStudy (anonymous):

y1 = 2 x1 = 1 m = -5/3 Therefore, y - 2 = (-5/3)(x - 1) This is in point-slope form